# Value Alignment of Social Media Ranking Algorithms

### v1

@@ A.2. Labeled Values
Appendix B Additional Study Details
We provide more information on the demographics of participants in both Study 1 and Study 2 as follows. In Study 1, our participants have a median age of $37.0$ years old. The sample is slightly skewed towards participants who identify as male ($52.5\%$). $45.4\%$ of participants identify as female, and $2.1\%$ as non-binary. Since we stratified our sampling of political leanings into two categories — Democrat and Republican or Independent, our participants are approximately balanced: $37.6\%$ are Democrats, $29.1\%$ are Independents, and $30.5\%$ are Republicans. With respect to race, $68.8\%$ identified as White, $13.5\%$ as Black, $5.7\%$ as Asian or Pacific Islander, $1.4\%$ as Indigenous / Native American, and $10.6\%$ as Multi-racial. Our participants are frequent users of Twitter with $62.7\%$ visiting the platform at least once a day.
Finally, we include information about the participants in our think-aloud study (Table 4). The median age of participants is $23.5$ years old. Nine of the participants identified as female, three as male, and one as non-binary. All but one of our participants use Twitter at least a few times a week.
@@ Appendix B Additional Study Details
B.1. Validating Labels
@@
We conduct an intrinsic evaluation of our LLM labeler on a dataset of 4,503 Twitter/X posts that have been labeled by at least three annotators. We report the disaggregated errors between any individual annotator and the remaining annotators as well as the difference between the LLM and the average annotations in Table 5.
B.2. Robustness Analysis

### v2

@@ A.2. Labeled Values
Appendix B Pilot Studies
We conducted three pilot studies: one for Study 1 and two for Study 2. For our pilot of Study 1, we recruited 30 participants from Prolific. To qualify for the study, participants must reside in the United States, be over 18 years old, have a Twitter account, and have over 250 submissions on Prolific with a $\geq 95\%$ acceptance rate. Participants received \$5.00 for completing the task (prorated from \$15.00 per hour). From the pilot, the mean recognizability rate was $75.0\%$.
For Study 2, we first ran a pilot where participants could change up to six sliders corresponding to different values. We recruited 150 participants from Prolific using the same screening criteria as for Study 1. Participants received \$5.75 for taking part in the study. From this pilot, we observed a large drop-off for participants who change more than one slider although the decrease in recognizability plateaus. Thus, for cost-saving measures, in the remaining pilots and in our final study, we allow participants to change up to five sliders or all nineteen. In the second pilot, we again recruited participants from Prolific using the same screening criteria. In total, we had 35 participants take part in the pilot; they received \$5 in compensation (prorated from \$15 per hour). In the study, participants are shown all 19 sliders but can adjust 1, 2, 3, 4, 5, or all 19 depending on which condition they have been randomly assigned. From our pilot results, we observe a mean recognizability of $70.5\%$ across conditions.
Appendix C Think-Aloud Study
To better understand how participants select value sliders, we conduct a think-aloud study where participants interact with the interface from Study 2. To note, these participants are distinct from those that took part in Studies 1 and 2.
C.1. Method
We provide details on how the think-aloud study was conducted and the analysis method. This study was approved by Stanford University’s Institutional Review Board.
C.1.1. Study Design
In the think-aloud study, participants completed the same protocol as described in Study 2 (Sec. 5) and were given access to all 19 sliders (i.e., the Full Sliders condition). We instructed the participants to think aloud as they adjusted their value sliders and reason about which feed they believed was value-ranked. Each session lasted approximately 35–60 minutes, and participants received a \$20 gift card via Tremendous as compensation. The think-aloud studies were video- and audio-recorded with participants’ consent.
C.1.2. Participants
In total, we recruited thirteen participants via university mailing lists. We include information about participant demographics in Table 4. The median age of participants is $23.5$ years old. Nine of the participants identified as female, three as male, and one as non-binary. All but one of our participants use Twitter at least a few times a week.
@@ C.1.2. Participants
C.1.3. Analysis
We analyzed the think-aloud sessions using an inductive, collaborative approach focused on identifying common patterns in how participants interacted with the slider interface and reasoned about value selections. We used Dovetail [^11] https://dovetail.com/ to transcribe the recordings from the think-aloud study. Then, two researchers independently read the transcripts, noting recurring behaviors and decision strategies. Then, we discussed these observations and iteratively developed a set of themes characterizing how participants interpreted the values, adjusted their sliders, and attempted to infer which feed was value-ranked. Using these collaboratively generated categories, both researchers revisited the transcripts and refined the themes through discussion. The resulting themes reflect consistent patterns in how participants understood the values, prioritized or deprioritized them, and navigated tensions between competing values.
C.2. Results
[…]
Figure 10. Overview of the study design flow with screenshots of the user interfaces.
Overview of the six-step study procedure The figure illustrates the full study workflow. First, participants provide access to their X/Twitter posts, which are collected for analysis. Second, they are redirected to the experimental survey platform. Third, participants complete a Portrait Value Questionnaire to assess personal values. Fourth, they adjust value sliders that dynamically reorder their social media posts based on selected values. Fifth, participants answer follow-up survey questions about task complexity and experience. Finally, they choose between different feed versions in a side-by-side comparison, selecting the one they prefer based on value alignment.
Appendix D Additional Study Details
First, we provide an overview of our study design in Fig. 10. Second, we provide more information on the demographics of participants in both Study 1 and Study 2 as follows. In Study 1, our participants have a median age of $37.0$ years old. The sample is slightly skewed towards participants who identify as male ($52.5\%$). $45.4\%$ of participants identify as female, and $2.1\%$ as non-binary. Since we stratified our sampling of political leanings into two categories — Democrat and Republican or Independent, our participants are approximately balanced: $37.6\%$ are Democrats, $29.1\%$ are Independents, and $30.5\%$ are Republicans. With respect to race, $68.8\%$ identified as White, $13.5\%$ as Black, $5.7\%$ as Asian or Pacific Islander, $1.4\%$ as Indigenous / Native American, and $10.6\%$ as Multi-racial. Our participants are frequent users of Twitter with $62.7\%$ visiting the platform at least once a day.
D.1. Validating Labels
@@
We conduct an intrinsic evaluation of our LLM labeler on a dataset of 4,503 Twitter/X posts from Epstein et al. (2025) that have been labeled by at least four annotators. The number of labels per post range from four to ten. Prior works have used human annotators to label expressions of Schwartz’s values in text content, including in written arguments (Kiesel et al., 2022), social scenarios (Qiu et al., 2022), and social media content (Epstein et al., 2025) Following these works, we filter to a set of posts that have been labeled by multiple human annotators to ensure higher confidence in the ground truth annotations, selecting a conservative threshold of posts with at least four annotators. We report the disaggregated errors between any individual annotator and the remaining annotators as well as the difference between the LLM and the average annotations in Table 5.
In addition, we also report the LLM-Consensus RMSE, or root mean-squared error, as well as the LLM-Consensus RMSE. Compared to MAE, RMSE is more sensitive to large errors, making it useful for identifying cases where the LLM’s predictions deviate substantially from the consensus. As reported in Table 6, the LLM-Consensus RMSE is $1.45$ whereas the Human-Consensus RMSE is $1.74$, indicating that the LLM performs as well, if not better, than the average annotator at predicting the consensus. Coupled with the LLM-Consensus MAE results, this indicates that our LLM labeler not only agrees closely with the consensus on average, but also avoids extreme mispredictions.
Value Human-Consensus RMSE LLM-Consensus RMSE
Self-directed thoughts $1.76$ $1.55$
Self-directed actions $1.80$ $1.60$
Stimulation $1.62$ $1.53$
Hedonism $1.87$ $1.50$
Achievement $1.65$ $1.58$
Dominance $1.28$ $1.09$
Resources $1.36$ $0.98$
Face $1.84$ $1.46$
Personal security $1.15$ $1.27$
Societal security $1.75$ $1.44$
Tradition $1.70$ $1.21$
Rule conformity $1.72$ $1.32$
Interpersonal conformity $1.58$ $1.43$
Humility $2.02$ $1.77$
Dependability $2.06$ $1.70$
Caring $2.19$ $1.77$
Universal concern $1.99$ $1.65$
Preservation of nature $1.41$ $1.08$
Tolerance $1.81$ $1.36$
Overall $1.74$ $1.45$
Table 6. For each value, we computed the root mean-squared error between any individual annotator and the remaining annotators (Human-Consensus RMSE) and the difference between the LLM and the average annotator (LLM-Consensus RMSE).
D.2. Robustness Analysis
D.2.1. Recognizability Measure
In Studies 1 and 2, we measure recognizability by asking the participant to select the feed they believe is more aligned to a single value or their selected values respectively. They are shown two side-by-side feeds: one feed is the value-aligned feed and the other is the engagement feed. As a robustness check, we replicated results from Study 2 when participants are shown one slider and all nineteen sliders in which participants are shown either their value-aligned feed or a random reshuffling of posts, both from the same batches of tweets. We recruited 80 participants in total; of these, 11 passed the attention checks, leaving us with 69 participants — 32 are allowed to adjust one slider and 37 are allowed to adjust all nineteen.
In line with results from Study 2, participants in the one slider condition had a mean recognizability of $71.1\%\pm 1.8$. For participants in the full sliders condition, mean recognizability was higher than in the original study at $69.4\%\pm 1.7$ compared to $61.0\%\pm 1.3$. It is possible that the engagement feed provides a more difficult baseline to compare against relative to a random reshuffling as there may be a small extent of user values embedded in the original ranking.
D.2.2. Sliders Changed versus Sliders Available